# FALCON: Fine-grained Activation Manipulation by Contrastive Orthogonal Unalignment for Large Language Model

### v1

@@ Abstract
Large language models have been widely applied, but can inadvertently encode sensitive or harmful information, raising significant safety concerns. Machine unlearning has emerged to alleviate this concern; however, existing training-time unlearning approaches, relying on coarse-grained loss combinations, have limitations in precisely separating knowledge and balancing removal effectiveness with model utility. In contrast, we propose Fine-grained Activation manipuLation by Contrastive Orthogonal uNalignment (FALCON), a novel representation-guided unlearning approach that leverages information-theoretic guidance for efficient parameter selection, employs contrastive mechanisms to enhance representation separation, and projects conflict gradients onto orthogonal subspaces to resolve conflicts between forgetting and retention objectives. Extensive experiments demonstrate that FALCON achieves superior unlearning effectiveness while maintaining model utility, exhibiting robust resistance against knowledge recovery attempts.
Machine Learning, ICML
@@ 1 Introduction
Recent advancements in generative AI (Achiam et al., 2023; Anil et al., 2023; Dubey et al., 2024), fueled by innovations in training techniques such as Parameter-Efficient Fine-Tuning (PEFT), have enabled large language models (LLMs) to efficiently internalize linguistic knowledge and excel in tasks from text generation to decision making (Hu et al., 2022; yang Liu et al., 2024). These models derive their power from massive, diverse corpora, but this dependency on large-scale datasets introduces significant risks: harmful, biased, or sensitive information can be inadvertently encoded and amplified, leading to ethical violations, regulatory noncompliance, and potential misuse (Hsu et al., 2024; Urman & Makhortykh, 2023; Jiao et al., 2024).
@@
Existing mitigation strategies, such as implementing guardrails (Dong et al., 2024) or training models with expertly curated datasets to refuse harmful queries (Ouyang et al., 2022), are computationally expensive and often inadequate against adversarial attacks (Xu et al., 2024; Yin et al., 2024). In contrast, while retraining an entire model on a cleaned dataset to eliminate harmful influences is theoretically feasible, it is prohibitively resource-intensive for modern LLMs (Jin et al., 2023). Additionally, adversaries can exploit PEFT to reintroduce such unwanted information, highlighting the urgent need for more effective and scalable solutions for publicly accessed LLMs (Qi et al., 2024).
@@
To solve harmful or sensitive information in machine learning models, Machine Unlearning (MU) has emerged as a promising solution, supported by growing regulations such as the “right to be forgotten” under the GDPR (Regulation, 2016; Ginart et al., 2019). It originally developed in the non-LLMs domain and has proven effective at removing specific data influences while preserving model performance (Bourtoule et al., 2021; Kim & Woo, 2022a; Liu et al., 2023). When transferred to maintain responsible LLMs, MU offers significant advantages, being far more computationally efficient than full retraining. Unlearned models also exhibit greater inherent safety, as they lack the undesired knowledge necessary for malicious behaviors (Hendrycks et al., 2021b; Li et al., 2024).
@@
Despite its potential, LLM unlearning still faces several fundamental issues: (1) existing approaches typically rely on empirical methods like grid search to identify intervention parameters, lacking an efficient and interpretable way to guide parameter selection within deep architectures, (2) current methods normally rely on coarse-grained manipulation (using simplistic loss combinations that induce random representation dispersion with uncontrolled gradient dynamics, struggling to balance knowledge removal and utility preservation) rather than fine-grained representation manipulation (achieving more effective knowledge separation through targeted representation modification and regulated gradient dynamics for reducing damage to model utility), and (3) knowledge recovery methods like jailbreaking attack can recover the undesired information from the unlearned model (Shaik et al., 2024).
To address the aforementioned issues of selective knowledge unlearning in LLMs, we propose Fine-grained Activation manipuLation by Contrastive Orthogonal uNalignment (FALCON), a representation-guided framework designed for precise knowledge removal while maintaining critical retention. FALCON leverages mutual information (MI) to evaluate dependencies between forget and retain data, enabling interpretable parameter selection and optimization guidance. For multi-domain knowledge unlearning, FALCON achieves fine-grained representation unalignment through two mechanisms: directional contrastive unalignment with singular value decomposition (SVD)-derived principal offset vectors and gradient orthogonal projection for precise control of knowledge separation granularity. Furthermore, our framework does not require full access to training data, making it practical for scenarios where such access is limited. FALCON further employs a second-order optimizer with approximate curvature information and can achieve efficient unlearning through single-layer parameter modification, ensuring both computational efficiency and scalability across diverse knowledge domains. Our contributions are as follows:
@@
  • We introduce information-theoretic metrics for quantifying knowledge entanglement, enabling principled parameter selection and providing theoretical insights into knowledge distribution across model architectures.
@@ 2 Related work
LLM unlearning refers to the selective removal of specific knowledge from large language models while preserving their overall functionality (Zhang et al., 2024a). Current approaches can be broadly categorized into training-time methods and inference-time methods (Barez et al., 2025). Among training-time approaches, which represent the mainstream methodology, two primary directions have emerged. The first direction focuses on gradient optimization, maximizing the loss function to suppress harmful knowledge through techniques such as gradient ascent (Yao et al., 2023; Jang et al., 2023) and reverse gradients (Eldan & Russinovich, 2023). The second direction emphasizes representation-guided adaptation, targeting specific intermediate hidden representations for modification (Li et al., 2024). While these aforementioned training-time methods achieve permanent unlearning by targeting specific layers and parameters, they currently rely heavily on coarse-grained loss combinations that struggle to disentangle deeply embedded knowledge representations flexibly.
Inference-time methods offer alternative approaches through task vectors and model editing. Task vector approaches address efficiency concerns through arithmetic operations on parameter-efficient modules, enabling lightweight unlearning under resource constraints (Ilharco et al., 2023; Zhang et al., 2023). Model editing usually modifies intermediate hidden states or logits to alter model behavior (Barez et al., 2025), such as contrastive decoding methods that suppress undesired content generation (Zhong et al., 2024). However, these methods’ dependence on modular arithmetic operations fundamentally limits their granularity in knowledge separation and constrains generalizability across diverse scenarios. Additionally, in-context unlearning has emerged as another inference-time approach, leveraging tailored prompts to dynamically suppress undesired outputs (Zheng et al., 2023; Pawelczyk et al., 2024). While flexible, this method’s effect remains inherently temporary as the undesired knowledge persists in the model’s representation space (Liu et al., 2024b).
@@ 3.1 Problem Setup
Figure 1: Schematic representation of the proposed FALCON method. The pipeline involves parameter selection based on mutual information, the construction of principal offset vectors, the application of contrastive mechanism on both datasets, and gradient conflict resolution to balance forgetting and retaining objectives.
The task of LLM unlearning involves selectively removing specific knowledge (forget set) from the model while retaining critical information (retain set). However, this process is complicated by the issue of knowledge entanglement, where representations of the forget and retain sets overlap significantly within the model’s parameters (Zhang et al., 2024b). This entanglement arises due to the distributed nature of knowledge across multiple layers and features, making it difficult to isolate knowledge for removal without affecting retained information.
To formalize the unlearning process, we adopt the general formulation proposed by Liu et al. (Liu et al., 2024b):
$\min_{\theta}\left\{\mathbb{E}_{(x,y_{f})\in\mathcal{D}_{\mathcal{F}}}\left[\mathcal{L}(y_{f}|x;\theta)\right]+\lambda\mathbb{E}_{(x,y)\in\mathcal{D}_{\mathcal{R}}}\left[\mathcal{L}(y|x;\theta)\right]\right\}$ (1)
@@
Despite the generality of above formulation, it does not explicitly quantify the representations of forgotten and retained knowledge. This lack of quantification poses challenges in precisely guiding the unlearning process (Qu et al., 2024). To address this, a principled metric is needed to evaluate and minimize knowledge entanglement, ensuring that unlearning primarily affects the forget set while minimizing interference with the retain set.
To address this, we introduce information-theoretic measures, specifically continuous entropy and mutual information, to quantify the dependency between the activations of the forget and retain sets. Let $\mathcal{F}$ and $\mathcal{R}$ represent the activations of the forget and retain sets at a specific layer of the model, respectively. The degree of knowledge entanglement between representations can be formulated as the mutual information $I(\mathcal{F};\mathcal{R})$:
$I(\mathcal{F};\mathcal{R})=H(\mathcal{F})+H(\mathcal{R})-H(\mathcal{F},\mathcal{R})$ (2)
where $H(\mathcal{F})$ and $H(\mathcal{R})$ are the continuous entropies of the activations $\mathcal{F}$ and $\mathcal{R}$, and $H(\mathcal{F},\mathcal{R})$ denotes their joint entropy. These measures provide a systematic approach to identify layers with minimal entanglement and guiding the optimization process for targeted and effective unlearning. The details of these metrics are shown in Appendix C.
@@ 3.2 LLM unlearning with MI Guidance
To quantify knowledge entanglement during machine unlearning, we use MI to measure the dependency between the activations of the forget set $\mathcal{F}^{(l)}$ and the retain set $\mathcal{R}^{(l)}$ at each layer $l$. The MI $I(\mathcal{F}^{(l)};\mathcal{R}^{(l)})$ serves as an indicator to guide the unlearning process by minimizing entanglement between $\mathcal{F}^{(l)}$ and $\mathcal{R}^{(l)}$.
To minimize the entanglement between the forget and retain sets’ representations, we formulate the layer selection as:
$l^{*}=\arg\min_{l}I(\mathcal{F}^{(l)};\mathcal{R}^{(l)})$ (3)
@@ 4 Methodology
To address the challenges of more thorough selective multi-domain knowledge unlearning and enhanced robustness against knowledge recovery in LLMs, we propose FALCON shown in Figure 1, a framework that advances both precision and effective in knowledge manipulation. Unlike prior approaches that rely on coarse-grained loss combinations, FALCON introduces three key mechanisms: (1) mutual information-based guidance to identify parameters where knowledge representations are least entangled, enabling interpretable parameter selection; (2) contrastive mechanism with enhanced representation separation to achieve fine-grained knowledge manipulation while ensuring robust resistance against knowledge recovery attempts; and (3) gradient orthogonal projection to resolve optimization conflicts and ensure training stability. This holistic design enables precise, interpretable, and robust knowledge unlearning in LLMs, transcending traditional loss-combination methods.
@@ 4.1 Information-Theoretic Guidance for Unlearning
In this paper, we utilize a principled approach to selective multi-domain knowledge unlearning in LLMs through mutual information. MI provides a natural measure of representational entanglement between the forget and retain datasets across model layers. By identifying layers that minimize MI, we can target unlearning interventions where forget and retain representations exhibit minimal overlap, thus preserving desired knowledge while selectively removing unwanted information.
@@
$I^{(l)}=\sum_{i=1}^{m}I(\mathcal{F}i^{(l)};\mathcal{R}^{(l)})+\eta\sum_{i=1}^{m}\sum_{j=i+1}^{m}I(\mathcal{F}_{i}^{(l)};\mathcal{F}_{j}^{(l)})$ (5)
where $m$ denotes the number of sub-domains in the forget set $\mathcal{F}$, and $\eta$ is a balancing coefficient that controls the relative importance of inter-domain dependencies.
For each layer $l$, since the activations are high-dimensional and continuous, direct entropy calculation is infeasible (Tsur et al., 2024). Instead, we utilize Kernel Density Estimation (KDE) to approximate the underlying global data distribution, estimating continuous entropy in activation space as defined in Appendix C (Walters-Williams & Li, 2009).
We implement KDE using a Multivariate Gaussian Kernel, which provides a smooth density estimation suitable for high-dimensional data. The estimated probability density function for activations $\mathcal{A}$ is given by:
$p(a)=\frac{1}{Nh}\sum_{n=1}^{N}K\left(\frac{a-a_{n}}{h}\right)$ (6)
where $a\in\mathbb{R}^{d}$ represents a single sample from the activations $\mathcal{A}$, including $\mathcal{F}$ and $\mathcal{R}$, with $d$ denoting the feature dimensionality of the activations, $N$ as the number of samples, $K(\cdot)$ represents the kernel function and $h$ as the adaptive bandwidth calculated using Scott’s rule (Scott, 2015), is defined as $h=\sigma N^{-\frac{1}{d+4}}$, which is particularly suitable for high-dimensional data due to its adjustment based on dimensionality. In this formula, $\sigma$ is the standard deviation of the data. This adaptive bandwidth selection can effectively balance bias and variance, ensuring robust density estimation for diverse activation distributions (Belhaj, 2024). Furthermore, to mitigate the curse of dimensionality, we apply Principal Component Analysis (PCA) to reduce the dimensions of the activations before performing KDE (Altman & Krzywinski, 2018). The number of components is chosen to retain at least 95% of the variance in the activation data, ensuring minimal information loss while significantly lowering computational complexity.
@@
$l^{*}=\arg\min_{l}\tilde{I}^{(l)}$ (7)
By identifying the layer with the lowest MI, we locate the model region where the forget and retain datasets are least entangled, minimizing the overlap between the two types of knowledge. Concurrently, this layer exhibits higher entanglement among sub-domains within the forget set, enabling efficient updates to shared representations across forgettable sub-domains. This dual property makes the layer the optimal focus for unlearning. During this process, its parameters are selectively optimized to remove undesired knowledge while preserving critical information, ensuring both precision in targeting the forget set and in retaining essential knowledge for downstream tasks.

### v2

@@ Abstract
Large language models have been widely applied, but can inadvertently encode sensitive or harmful information, raising significant safety concerns. Machine unlearning has emerged to alleviate this concern; however, existing training-time unlearning approaches, relying on coarse-grained loss combinations, have limitations in precisely separating knowledge and balancing removal effectiveness with model utility. In contrast, we propose Fine-grained Activation manipuLation by Contrastive Orthogonal uNalignment (FALCON), a novel representation-guided unlearning approach that leverages information-theoretic guidance for efficient parameter selection, employs contrastive mechanisms to enhance representation separation, and projects conflict gradients onto orthogonal subspaces to resolve conflicts between forgetting and retention objectives. Extensive experiments demonstrate that FALCON achieves superior unlearning effectiveness while maintaining model utility, exhibiting robust resistance against knowledge recovery attempts. Our implementation is available at: https://github.com/CharlesJW222/FALCON/tree/main
footnotetext: Corresponding authors: {yi.dong, xiaowei.huang}@liverpool.ac.uk
@@ 1 Introduction
Recent advancements in generative AI achiam2023gpt ; dubey2024llama , powered by Parameter-Efficient Fine-Tuning (PEFT) techniques, have enabled LLMs to internalize linguistic knowledge and excel across diverse tasks ao2025safepruninglorarobust ; hu2022lora . While these models gain their capabilities from massive datasets, this reliance on large-scale corpora creates significant risks: harmful, biased, or sensitive information can become encoded and amplified, resulting in ethical violations, regulatory noncompliance, and potential misuse hsu2024safe ; urman2023silence ; jiao2024navigating .
@@
Existing mitigation strategies, such as LLM guardrails dong2024building or training models with expertly curated datasets to refuse harmful queries NEURIPS2022_b1efde53 , are computationally expensive and often inadequate against adversarial attacks yin2024vqattack . In contrast, while retraining an entire model on a cleaned dataset to eliminate harmful impacts is theoretically feasible, it is prohibitively resource-intensive for modern LLMs 10431584 . Additionally, adversaries can exploit PEFT to reintroduce such unwanted information, highlighting the urgent need for more effective and scalable solutions for publicly accessed LLMs qi2024finetuning .
@@
To solve harmful or sensitive information in machine learning models, Machine Unlearning (MU) has emerged as a promising solution, supported by growing regulations such as the “right to be forgotten" under the GDPR regulation2016regulation . It commonly developed in the non-LLMs domain and has proven effective at removing specific data influences while preserving model performance liu2023muter ; cha2024learning ; wu2025unlearning . When transferred to maintain responsible LLMs, MU offers significant advantages, being far more computationally efficient than full retraining. Unlearned models also exhibit greater inherent safety, as they lack the undesired knowledge necessary for malicious behaviors hendrycks2021unsolved ; li2024the .
@@
Despite its potential, LLM unlearning still faces several fundamental issues: (I1) existing approaches typically rely on empirical methods like grid search to identify intervention parameters, lacking efficient and interpretable guidance within deeper LLM architectures, (I2) current methods normally rely on coarse-grained manipulation (using simplistic loss combinations that induce random representation dispersion with uncontrolled gradient dynamics, struggling to balance knowledge removal and utility preservation) rather than fine-grained representation manipulation (achieving more effective knowledge separation through targeted representation modification and regulated gradient dynamics for reducing damage to model utility), and (I3) knowledge recovery methods such as jailbreaking attack can recover the undesired information from the unlearned model 10750906 .
Figure 1: Schematic overview of FALCON. The pipeline comprises three stages: parameter selection based on mutual information (Step 1); contrastive orthogonal unalignment, which consists of contrastive mechanism on both forgetting and retention datasets (Step 2.1) and orthogonal gradient conflict resolution (Step 2.2); and model unlearning guided by these components (Step 3).
To address the aforementioned issues of selective knowledge unlearning in LLMs, we propose Fine-grained Activation manipuLation by Contrastive Orthogonal uNalignment (FALCON), a representation-guided framework for targeted knowledge removal with minimal impact on general capabilities. For I1, FALCON uses mutual information (MI) as an auxiliary signal to assess dependencies between forget and retain data, based on which it introduces two core mechanisms for fine-grained disentanglement and unlearning (Step 1). To tackle I2, FALCON utilizes singular value decomposition (SVD) to identify principal directions in activation space to steer representations along axes misaligned with forgettable knowledge, enabling more thorough removal (Step 2.1). Meanwhile, FALCON uses a gradient orthogonal projection strategy, which constrains updates away from retention-sensitive directions, reducing interference with preserved content (Step 2.2). These mechanisms enable precise unlearning with limited data access and remain effective even under single-layer interventions. Afterwards, the projected gradients are used to update the model parameters (Step 3). For I3, we provide comprehensive empirical evidence and analysis in Section 5.3 and Appendix E.6 to support our claims. Our contributions are as follows:
@@
  • We introduce information-theoretic metrics for quantifying knowledge entanglement, enabling principled parameter selection and providing empirical insights into knowledge distribution across model architectures.
@@ 2 Related work
LLM unlearning refers to the selective removal of specific knowledge from large language models while preserving their overall functionality zhang2024right . Current approaches can be broadly categorized into training-time methods and inference-time methods barez2025openproblemsmachineunlearning . Among training-time approaches, which represent the mainstream methodology, two primary directions have emerged. The first direction focuses on gradient optimization yao2023large ; jang-etal-2023-knowledge ; eldan2023s ; zhang2024negative ; fan2024simplicity , which suppresses harmful knowledge through loss-driven techniques but often causes catastrophic forgetting and instability when distributions are highly similar or lack fine-grained knowledge localization. The second direction emphasizes representation-guided adaptation, targeting intermediate hidden representations for modification li2024the ; zou2024improving ; rosati2024representation , but relying on empirical layer selection and lacking targeted separation mechanisms. While these aforementioned training-time methods achieve permanent unlearning by targeting specific layers and parameters, they currently rely heavily on coarse-grained loss combinations that struggle to disentangle deeply embedded knowledge representations flexibly jia2024wagle .
Inference-time methods offer alternative approaches like task vectors and model editing. Task vector approaches address efficiency concerns through arithmetic operations on parameter-efficient modules, enabling lightweight unlearning under resource constraints ilharco2023editing ; zhang2023composing , but oversimplify knowledge structure through linear assumptions that fail to capture complex knowledge entanglement. In contrast, model editing usually modifies intermediate hidden states or logits to alter model behavior barez2025openproblemsmachineunlearning ; ji2024reversing ; dong2024undial ; huang2024offset , such as contrastive decoding methods that prevent inappropriate responses zhong2024rose . Moreover, ECO liu2024large has also demonstrated promising performance, though it functions more as a guardrail’s definition for filtering sensitive content dong2024safeguardinglargelanguagemodels ; hu2025trustorientedadaptiveguardrailslarge , rather than directly serving as an unlearning algorithm [^1] Further discussion on ECO is shown in Appendix. F.3 ucki2024an . However, these methods’ dependence on modular arithmetic operations fundamentally limits their granularity in knowledge separation and constrains generalizability across diverse scenarios. Additionally, in-context unlearning has emerged as another inference-time approach, leveraging tailored prompts to dynamically suppress undesired outputs zheng2023can ; pawelczyk2024incontext . While flexible, this method’s effect remains inherently temporary as the undesired knowledge persists in the model’s representation space liu2024rethinking .
@@ 3.1 Problem Setup
The task of LLM unlearning involves selectively removing specific knowledge (forget set) from the model while retaining critical information (retain set). However, this process is complicated by the issue of knowledge entanglement, where representations of the forget and retain sets overlap significantly within the model’s parameters zhang2024comprehensive . This entanglement arises due to the distributed nature of knowledge across multiple layers and features, making it difficult to isolate knowledge for removal without affecting retained information. To formalize the unlearning process, we adopt the general formulation proposed by Liu et al. liu2024rethinking :
$\small\min_{\theta}\left\{\mathbb{E}_{(x,y_{f})\in\mathcal{D}_{\mathcal{F}}}\left[\mathcal{L}(y_{f}|x;\theta)\right]+\lambda\mathbb{E}_{(x,y)\in\mathcal{D}_{\mathcal{R}}}\left[\mathcal{L}(y|x;\theta)\right]\right\}$ (1)
@@
Despite the generality of above formulation, it does not explicitly quantify the representations of forgotten and retained knowledge. This lack of quantification poses challenges in precisely guiding the unlearning process qu2024frontier . To address this, a principled metric is needed to evaluate and minimize knowledge entanglement, ensuring that unlearning primarily affects the forget set while minimizing interference with the retain set. Consequently, we introduce information-theoretic measures, specifically continuous entropy and mutual information, to quantify the dependency between the activations of the forget and retain sets. Let $\mathcal{F}$ and $\mathcal{R}$ represent the activations of the forget and retain sets at a specific layer of the model, respectively. The degree of knowledge entanglement between representations can be formulated as the MI $I(\mathcal{F};\mathcal{R})$:
$\small I(\mathcal{F};\mathcal{R})=H(\mathcal{F})+H(\mathcal{R})-H(\mathcal{F},\mathcal{R})$ (2)
where $H(\mathcal{F})$ and $H(\mathcal{R})$ are the continuous entropies of the activations $\mathcal{F}$ and $\mathcal{R}$, and $H(\mathcal{F},\mathcal{R})$ denotes their joint entropy. These measures provide a systematic approach to identify parameters with minimal entanglement and guide the LLM unlearning process. The details of these metrics are shown in Appendix C.
@@ 3.2 LLM unlearning with MI Guidance
To quantify knowledge entanglement during machine unlearning, we use MI to measure the dependency between the activations of the forget set $\mathcal{F}^{(l)}$ and the retain set $\mathcal{R}^{(l)}$ at each layer $l$. The MI $I(\mathcal{F}^{(l)};\mathcal{R}^{(l)})$ serves as an indicator to guide the unlearning process by minimizing entanglement between $\mathcal{F}^{(l)}$ and $\mathcal{R}^{(l)}$. To minimize the entanglement between the forget and retain sets’ representations, we formulate the parameter selection for specific LLM layers as:
$\small l^{*}=\arg\min_{l}I(\mathcal{F}^{(l)};\mathcal{R}^{(l)})$ (3)
@@ 4 Methodology
To address the challenges of more thorough selective multi-domain knowledge unlearning and enhanced robustness against knowledge recovery in LLMs, we propose FALCON shown in Figure 1 and Appendix. D.1, a framework that advances both precision and effectiveness in knowledge manipulation. Unlike prior approaches that rely on coarse-grained loss combinations, FALCON introduces three key mechanisms: (1) mutual information-based guidance to identify parameters where knowledge representations are least entangled, enabling interpretable parameter selection; (2) contrastive mechanism with enhanced representation separation to achieve fine-grained knowledge manipulation while ensuring robust resistance against knowledge recovery attempts; and (3) gradient orthogonal projection to resolve optimization conflicts and ensure training stability. This holistic design enables precise, interpretable, and robust knowledge unlearning in LLMs, transcending traditional loss-combination methods.
@@ 4.1 Information-Theoretic Guidance for Unlearning
In this paper, we utilize a principled approach to selective multi-domain knowledge unlearning in LLMs through mutual information. MI provides a natural measure of representational entanglement between the forget and retain datasets across model layers. By identifying parameters that minimize MI, we can target unlearning interventions where forget and retain representations exhibit minimal overlap, thus preserving desired knowledge while selectively removing unwanted information.
@@
$\small I^{(l)}=\sum_{i=1}^{m}I(\mathcal{F}i^{(l)};\mathcal{R}^{(l)})+\eta\sum_{i=1}^{m}\sum_{j=i+1}^{m}I(\mathcal{F}_{i}^{(l)};\mathcal{F}_{j}^{(l)})$ (5)
where $m$ denotes the number of sub-domains in the forget set $\mathcal{F}$, and $\eta$ is a balancing coefficient that controls the relative importance of inter-domain dependencies. For each layer $l$, since the activations are high-dimensional and continuous, direct entropy calculation is infeasible tsur2024max . Instead, we utilize Kernel Density Estimation (KDE) to approximate the underlying global data distribution, estimating continuous entropy in activation space as defined in Appendix C walters2009estimation . Specifically, we use a multivariate Gaussian kernel, which offers a smooth and flexible density estimation well-suited to high-dimensional data. The estimated probability density function for activations $\mathcal{A}$ is given by:
$\small p(a)=\frac{1}{Nh}\sum_{n=1}^{N}K\left(\frac{a-a_{n}}{h}\right)$ (6)
where $a\in\mathbb{R}^{d}$ represents a single sample from the activations $\mathcal{A}$, including $\mathcal{F}$ and $\mathcal{R}$, with $d$ denoting the feature dimensionality of the activations, $N$ as the number of samples, $K(\cdot)$ represents the kernel function and $h$ as the adaptive bandwidth calculated using Scott’s rule scott2015multivariate , defined as $h=\sigma N^{-\frac{1}{d+4}}$, which is particularly suitable for high-dimensional data due to its dimensionality-based adjustment. Here, $\sigma$ is the standard deviation of the data. This adaptive bandwidth selection effectively balances bias and variance, ensuring robust density estimation for diverse activation distributions belhaj2024modified . To mitigate the curse of dimensionality, we apply Principal Component Analysis (PCA), which has been widely adopted across various domains in prior work kleindessner2023efficient ; qiu2024spectral ; shao-etal-2023-gold to reduce activation dimensions before performing KDE altman2018curse , retaining at least 95% of variance to ensure minimal information loss while significantly lowering computational complexity.
@@
$\small l^{*}=\arg\min_{l}\tilde{I}^{(l)}$ (7)
By identifying the layer with the lowest MI, we locate the model region where the forget and retain datasets are least entangled, minimizing the overlap between the two types of knowledge. Concurrently, this layer exhibits higher entanglement among sub-domains within the forget set, enabling efficient updates to shared representations across forgettable sub-domains. This dual property makes the layer an optimal target for unlearning, where parameters with minimal mutual interference are prioritized to remove undesired knowledge while more easily preserving essential and generalizable knowledge for downstream tasks.